### v1

# Bitcoin Under Volatile Block Rewards: How Mempool Statistics Can Influence Bitcoin Mining

Abstract.
As Bitcoin experiences more halving events, the protocol reward converges to zero, making transaction fees the primary source of miner rewards. This shift in Bitcoin’s incentivization mechanism, which introduces volatility into block rewards, could lead to the emergence of new security threats or intensify existing ones. Previous security analyses of Bitcoin have either considered a fixed block reward model or a highly simplified volatile model, overlooking the complexities of Bitcoin’s mempool behavior.
In this paper, we present a reinforcement learning-based tool designed to analyze mining strategies under a more realistic volatile model. Our tool uses the Asynchronous Advantage Actor-Critic (A3C) algorithm to derive near-optimal mining strategies while interacting with an environment that models the complexity of the Bitcoin mempool. This tool enables the analysis of adversarial mining strategies, such as selfish mining and undercutting, both before and after difficulty adjustments, providing insights into the effects of mining attacks in both the short and long term.
Our analysis reveals that Bitcoin users’ trend of offering higher fees to speed up the inclusion of their transactions in the chain can incentivize payoff-maximizing miners to deviate from the honest strategy. In the fixed reward model, a disincentive for the selfish mining attack is the initial loss period of at least two weeks, during which the attack is not profitable. However, our analysis shows that once the protocol reward diminishes to zero in the future, or even currently on days when transaction fees are comparable to the protocol reward, mining pools might be incentivized to abandon honest mining to gain an immediate profit.
Bitcoin, Selfish mining, Undercutting, Reinforcement learning, A3C
1. Introduction
As of September 2024, Bitcoin (Nakamoto, 2008) holds a market capitalization of nearly 1.18 trillion dollars (mar, 2024), surpassing the annual Gross Domestic Product (GDP) of over 90% of countries (gdp, 2024). While Bitcoin has progressed without major security attacks thus far, its steady growth over the past decade and a half does not guarantee future security, highlighting the need to address potential vulnerabilities. Protecting Bitcoin’s economy requires ensuring the resilience of its incentivization mechanism. If consensus nodes find their rewards insufficient to cover costs, they may abandon the chain, risking a halt in blockchain operations. Bitcoin’s incentivization comes from two sources: the protocol reward for mining a block and the transaction fee paid by users. In Bitcoin’s early years, the protocol reward significantly outweighed transaction fees. However, due to the halving mechanism, which reduces the protocol reward every four years, transaction fees are expected to match, and eventually replace, the protocol reward as the sole source of incentivization. In fact, there have already been periods, such as in December 2023 and April 20, 2024, when transaction fees equaled or even exceeded the protocol reward (exp, 2024).
As the balance between transaction fees and the protocol reward changes over time, the incentivization dynamics of Bitcoin change as well, potentially leading to the emergence of new security threats or intensifying existing ones. One consequence of transitioning to a transaction-fee-driven era is that the total reward per block can no longer be considered fixed. Block rewards become volatile, as different blocks include varying transactions with different levels of transaction fees. This volatility in block rewards alters the profitability of various mining strategies. For example, under the fixed reward model, when only a single fork exists in the network, miners are incentivized to mine on the longest chain (Sapirshtein et al., 2017). Since the block reward is fixed, there is no benefit in taking the risk of orphaning previous blocks. However, in a volatile block reward model, a miner might find it more profitable to attempt to orphan a previously mined block with a significantly high transaction fee to steal those transactions (Bar-Zur et al., 2022; Carlsten et al., 2016). Additionally, a miner may choose to mine a secret fork of blocks to prevent others from building on that fork, allowing the miner to include all available transactions from the mempool in its secret fork. These examples highlight the need to revisit the security analyses conducted under the fixed reward model (Zur et al., 2020; Sapirshtein et al., 2017; Grunspan and Pérez-Marco, 2018) to better understand Bitcoin’s future security in the context of volatile block rewards.
Some research has analyzed Bitcoin’s security under the volatile block reward model. Carlsten et al. (Carlsten et al., 2016) raised concerns about Bitcoin security as the protocol reward diminishes to zero, introducing the undercutting attack and showing how selfish mining could become more threatening in a transaction-fee-driven era. While the paper highlights Bitcoin’s instability under the volatile reward model, the analysis relies on simplified assumptions, such as unlimited block space, allowing blocks to collect all available transaction fees. Additionally, the analysis is based on a basic selfish mining strategy, overlooking the optimal strategy and other actions that adversarial miners might use in the volatile block reward model. To improve mining analysis in the transaction-fee setting, the authors in (Bar-Zur et al., 2022) introduced the WeRLman framework, which combines deep reinforcement learning with a Monte Carlo Tree Search (MCTS) tool, inspired by the AlphaGo Zero implementation (Silver et al., 2017). They assumed occasional whale transactions with significantly high fees that can result in volatility of block rewards. The study showed that as Bitcoin undergoes more halvings, the minimum hash-rate needed for profitable deviation from honest mining decreases. While the WeRLman paper made progress in analyzing Bitcoin security, its mempool implementation and treatment of transaction fees do not accurately reflect real-world conditions. In WeRLman, a block can include either no whale transactions or just one, resulting in rewards of either 1 or $1+\texttt{fee}$, where fee is the whale transaction fee. Therefore, the WeRLman framework simplifies the sophisticated mempool environment to only the count of whale transactions, all with the same fee. Moreover, WeRLman is designed to obtain a near-optimal strategy only after adjusting the mining difficulty. However, in practice, if a miner deviates from the honest strategy, it takes an epoch of at least two weeks for the difficulty to be adjusted, a relatively long period that WeRLman overlooks in its analysis.
In this paper, we enhance the analysis of mining strategies in Bitcoin by incorporating a more realistic volatile block reward model that obtains the near-optimal strategies both before and after the difficulty adjustment. We introduce two methods for evaluating Bitcoin security: the simplified method and the enhanced method. In the simplified method, we analyze historical Bitcoin block data over various periods to establish the relationship between the time required to mine a block, referred to as block generation time, and the total amount of transaction fees included in the block. This relationship is used to develop a reward function that approximates the behavior of the mempool under the volatile reward model. To evaluate mining strategies before difficulty adjustment within this simplified volatile reward framework, we employ a Markov Decision Process (MDP)-based tool. Because the MDP can handle a limited state space, our tool calculates a lower bound for miners’ profitability before a difficulty adjustment. This lower bound is particularly useful for determining the threshold mining parameters at which deviating from honest mining becomes profitable.
In the enhanced method, we improve both the implementation of mempool behavior and the training approach. Specifically, we apply the Asynchronous Advantage Actor-Critic (A3C) reinforcement learning algorithm (Mnih et al., 2016) to derive a near-optimal mining strategy. In this algorithm, multiple agents (representing a payoff-maximizing miner) interact with their own instance of the environment (the Bitcoin blockchain) to simultaneously train a global policy. To better model the Bitcoin mempool in the A3C environment, we analyze historical Bitcoin mempool data to estimate the transaction fee arrival patterns. In our implementation, transactions are categorized into different groups based on the fee paid per unit of weight. For each category, we estimate how the total transaction weight grows over time, incorporating these estimates into our environment. Our A3C implementation considers a larger set of actions and states compared to models designed for the fixed reward system, enabling it to obtain near-optimal policies both before and after difficulty adjustments under the volatile reward model.
Our analysis shows that when transaction rewards become comparable to protocol rewards—whether in the future with a significantly low protocol reward or currently on days with relatively high transaction fees—miners do not need to wait for infrequent whale transactions to be incentivized to deviate from the honest strategy. The inherent competition among transactions to be included in a block creates an opportunity for payoff-maximizing miners to earn more than their fair share of rewards by deviating from the honest strategy. As more time passes after the last block is mined on the network, new transactions tend to raise their fees to improve their chances of being included in the next block. In this situation, if a malicious miner removes other miners’ blocks from the canonical chain, it can increase the generation time of its own blocks. This extended block generation time allows the miner to include transactions with higher fees in its blocks.
Another interesting observation is that under the volatile reward model, a deviant strategy, such as a selfish mining attack, can become profitable immediately after being launched without needing a mining difficulty adjustment. As discussed in the literature (Grunspan and Pérez-Marco, 2018; Negy et al., 2020; Grunspan and Pérez-Marco, 2023; Sarenche et al., 2024b), under the fixed reward model, a selfish miner incurs losses during the initial epoch of attack when the mining difficulty has not yet adjusted, the duration of which is at least two weeks. This initial loss period is considered one of the main reasons such attacks have not been observed in the Bitcoin network. However, under the volatile reward model, this loss period is shortened or may disappear entirely, making selfish mining a more significant threat in this era.
Figure 1. Immediate profitability of selfish mining in the transaction-fee era.
Once the protocol reward reaches zero, a threatening pattern that can emerge in the Bitcoin mempool occurs when the mempool is relatively empty, and blocks primarily collect transactions that arrive during their generation time. This mempool pattern, observable in Bitcoin’s mempool during September 2024, is susceptible to deviant strategies such as selfish mining and undercutting attacks. The graph depicted in Figure 1 represents a lower bound for the percentage increase in a miner’s time-averaged profit before a difficulty adjustment, on a day when the protocol reward is zero, and the same mempool pattern occurs as on September 5, 2024. This time-averaged profit percentage increase is calculated for a miner with a mining share equal to $1/3$, as a function of its communication capability, which denotes the ratio of miners receiving adversarial blocks earlier in the event of a block race. The normal communication capability is assumed to be $50\%$. However, as demonstrated in (Göbel et al., 2016), for a well-distributed mining pool, its connectivity ratio can increase substantially. As illustrated in Figure 1, the initial loss period of selfish mining can be entirely eliminated for malicious miners in the transaction-fee era.
To summarize, our key contributions are:
  • Discussion on how a volatile block reward model can affect selfish mining profitability, both before and after the difficulty adjustment, presented in Section 3.
  • Implementation of an MDP-based tool to calculate a lower bound on selfish mining profitability under a simplified mempool implementation, presented in Section 4.
  • Implementation of an A3C-based tool to obtain a near-optimal mining strategy under a realistic mempool model, presented in Section 5.
To present the results in this paper, we have run our implementation using real-world mempool statistics (mem, 2024) from different periods to demonstrate the threat of adversarial deviation from the honest strategy if the same mempool patterns reoccur.
2. System Model and Definitions
The honest strategy in Bitcoin protocol is defined as follows:
Definition 2.1 (Honest strategy).
A mining node should always mine on top of the longest chain  [^1] In Proof-of-Work blockchains, the longest chain is a chain with the highest accumulated difficulty. available in its view. Once a new block is mined, the mining node should immediately publish the block to all the other mining nodes.
The mining nodes in our scheme are categorized into two groups:
  • Honest mining nodes: These nodes, which are denoted by $\mathcal{H}$, follow the honest strategy defined in Definition 2.1.
  • Adversarial mining nodes: In our paper, we assume that the adversary, denoted by $\mathcal{A}$, controls a subset of mining nodes referred to as adversarial nodes. These adversarial nodes can arbitrarily deviate from the honest strategy. Namely, they can choose any fork rather than the longest chain to mine on top of it or withhold the mined blocks.
A brief discussion of mining attacks that an adversary can conduct in our model, including selfish mining and undercutting, is presented in Appendix A. Depending on the context, we may refer to adversarial nodes as an attacker, a selfish miner, or an undercutter. Although the longest chain fork choice rule outputs a deterministic chain in most cases, situations may arise in which there are multiple forks with the same height in the view of a mining node (the same-height fork race), making the chain selection somewhat subjective. Based on how mining nodes select the winning chain in the presence of multiple same-height forks, the honest nodes are divided into $2$ sub-groups:
  • Altruistic mining nodes: These nodes always choose a fork they have heard about earlier in the case of a same-height fork race.
  • Petty-compliant mining nodes: These nodes choose a fork that leads to a greater profit for them in the case of a same-height fork race.
According to the definition above, in a same-height fork race, petty-compliant nodes choose the fork with the lower total transaction fees included. The rationale behind this choice is that the fork with fewer transaction fees leaves more fees in the transaction pool, allowing the next block mined on this fork to include a higher amount of fees. Assuming all mining nodes share the same view of the transaction pool, in a same-height fork race, the winning chain for a petty-compliant node is the one whose tip block (the highest block) was mined earlier and, as a result, includes lower fees.
Another factor, besides mining share, that can affect the adversary’s profitability in following a deviant strategy instead of honest mining is its connectivity to other mining nodes. We define adversarial communication capability as follows:
Definition 2.2 (Adversarial communication capability).
The adversarial communication capability is defined as the ratio of altruistic nodes that receive the adversarial fork earlier than the honest fork in the case of a same-height fork race.
The adversarial communication capability specifies the ratio of altruistic nodes that select the adversarial chain as the winning one in a same-height fork race. Let $\alpha$ denote the mining power share corresponding to the adversary, $\beta$ denote the ratio of petty-compliant mining nodes to the total honest mining nodes, and $\gamma$ denote the adversarial communication capability. This implies that the adversarial, petty-compliant, and altruistic mining shares are equal to $\alpha$, $\beta(1-\alpha)$, and $(1-\beta)(1-\alpha)$, respectively. Under the same-height fork race, the ratio $\gamma(1-\beta)(1-\alpha)$ of mining share (referring to the altruistic nodes that receive the adversarial chain earlier) always mines on top of the adversarial chain.
In our model, time is assumed to be divided into rounds, denoted by $r$. In each round, a mining node can calculate multiple mining (hash) queries, the number of which is proportional to its mining share. A time step, denoted by $t$, represents the set of consecutive rounds after which a block is mined in the system. At each time step $t$, a block is generated by either the adversary or the honest miners. However, depending on the adversary’s strategy, the block generated at time step $t$ might not be immediately added to the canonical chain. Instead, it could be either added to the canonical chain or eventually be identified as an orphan block (excluded from the canonical chain) at a future time step.
Definition 2.3 (Block ratio).
The block ratio of adversary $\mathcal{A}$ following strategy $\pi$ is defined as follows:
(1) $\begin{split}&\texttt{BR}^{\mathcal{A}}_{T}(\pi)=\frac{\sum_{t=0}^{T}{N_{\mathcal{A}}(t;\pi)}}{\sum_{t=0}^{T}{N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)}},\;\;\;\;\text{and}\;\;\;\;\\ &\texttt{BR}^{\mathcal{A}}(\pi)=\lim_{T\to\infty}{\texttt{BR}^{\mathcal{A}}_{T}(\pi)},\end{split}$
where $N_{\mathcal{A}}(t;\pi)$ and $N_{\mathcal{H}}(t;\pi)$ denote the number of adversarial blocks and the number of honest blocks, respectively, added to the canonical chain at time step $t$ under policy $\pi$.
Definition 2.4 (Time-averaged profit).
The time-averaged profit (per-round profit) of adversary $\mathcal{A}$ following strategy $\pi$ is defined as follows:
(2) $\begin{split}&\texttt{Profit}^{\mathcal{A}}_{T}(\pi)={\frac{\sum_{t=1}^{T}{R_{\mathcal{A}}(t;\pi)-C_{\mathcal{A}}(t;\pi)}}{T}},\;\;\;\;\text{and}\;\;\;\;\\ &\texttt{Profit}^{\mathcal{A}}(\pi)=\lim_{T\to\infty}{\texttt{Profit}^{\mathcal{A}}_{T}(\pi)}\enspace,\end{split}$
where $R_{\mathcal{A}}(t;\pi)$ and $C_{\mathcal{A}}(t;\pi)$ denote the revenue achieved and the mining cost incurred by adversary $\mathcal{A}$ at time step $t$ under policy $\pi$, respectively.
The revenue a miner receives at a given time step equals the total rewards from its blocks added to the canonical chain at that time. Each block reward consists of the protocol reward and the transaction fees included in the block.
3. Selfish Mining Profitability
This section provides an intuitive discussion on selfish mining profitability in the context of Bitcoin. We explain why selfish mining can be a profitable deviant strategy and explore the effect of the difficulty adjustment mechanism (DAM) on its profitability. We also discuss why selfish mining always incurs an initial period of loss in the fixed reward block model. Additionally, we examine how transitioning to the volatile block reward model can influence selfish mining profitability both before and after the difficulty adjustment mechanism. This discussion can aid in better understanding the results obtained through implementation in Sections 4 and 5.
3.1. Non-profitability of Pre-DAM Selfish Mining Under the Fixed Block Reward Model
In Bitcoin, a difficulty epoch is defined as the period during which 2016 blocks are appended to the canonical chain. At the end of each epoch, the Difficulty Adjustment Mechanism (DAM) recalibrates the difficulty target for the subsequent epoch based on the hash power estimated from the previous epoch. Under a fixed block reward model, selfish mining cannot be more profitable than honest mining during the first epoch and before the DAM is applied (Grunspan and Pérez-Marco, 2018). To illustrate this, it is essential to compare the adversary’s time-averaged profit from both honest and selfish mining prior to the difficulty adjustment. Let $e$ denote an epoch and $\lambda$ denote the mining rate, assuming all mining nodes, including the adversary, mine honestly. In this case, $\frac{1}{\lambda}$ represents the average number of rounds required to add a new block to the canonical chain. We define the adversary’s total block generation rate as the average number of rounds in which a new adversarial block is mined. We also define the adversary’s canonical block generation rate as the average number of rounds in which a new adversarial block is added to the canonical chain. Under honest mining, both the total and canonical block generation rates are the same. In the fixed-block reward model, we assume the reward for each block in the canonical chain is constant, denoted by $R$. We ignore the mining cost for simplicity, as it is the same under both honest and selfish mining.
If the adversary $\mathcal{A}$, with a stake share of $\alpha$, mines honestly during epoch $e$, its canonical block generation rate would be equal to $\lambda\alpha$, resulting in a time-averaged profit of $\lambda\alpha R$. If the adversary $\mathcal{A}$ starts selfish mining at the beginning of epoch $e$, its canonical block generation rate would be less than or equal to that under honest mining ($\leq\lambda\alpha$). This is because the difficulty of epoch $e$ is already set before the start of the epoch, and the adversary’s strategy during the epoch cannot influence the epoch difficulty. Since both the epoch difficulty and the adversary’s mining share are fixed, the adversary’s total block generation rate remains the same under both honest and selfish mining. However, under selfish mining, some of the adversarial blocks may become orphaned, leading to a decrease in the adversary’s canonical block generation rate. Therefore, the time-averaged profit from selfish mining during epoch $e$ (before a DAM is applied) is $\leq\lambda\alpha R$. Intuitively, before a DAM is applied, the adversary cannot mine blocks faster, regardless of the strategy chosen (Sarenche et al., 2024b).
Note that while selfish mining may potentially increase the adversary’s block ratio during epoch $e$, this increase is due to orphaned honest blocks and does not reflect an increase in the adversary’s time-averaged profit.
3.2. Profitability of Long-Range Selfish Mining Under the Fixed Block Reward Model
Knowing that under the fixed block reward model, selfish mining cannot be profitable during the initial epoch, the adversary needs to sustain the selfish mining attack for a longer period to achieve profitability. In Bitcoin, DAM is responsible for adjusting the block generation rate to ensure that it takes, on average, 10 minutes to mine a new block. This implies that the ideal epoch duration is two weeks. Assume the adversary starts selfish mining at the beginning of epoch $e$. As a result of selfish mining, some of the blocks (both adversarial and honest) mined during epoch $e$ may become orphaned. The generation of these orphan blocks extends the duration of epoch $e$ because to add 2016 blocks to the canonical chain, more than 2016 blocks need to be mined. At the end of epoch $e$, which exceeds two weeks, the DAM decreases the mining difficulty for the following epoch to restore the canonical block generation rate to one block per 10 minutes. This reduction in difficulty leads to a higher block generation rate for the adversary, indicating that the adversary can mine new blocks in fewer rounds (Sarenche et al., 2024b).
This implies that under the fixed block reward model, selfish mining requires at least one difficulty epoch to become profitable. In practice, the attack duration must be extended further to ensure profitability. This is because the selfish miner incurs losses during the initial epoch due to orphaned blocks, necessitating a longer attack period to compensate for these losses. Let $\texttt{BR}^{\mathcal{A}}(\pi^{\texttt{selfish}})$ denote the adversarial block ratio under selfish mining, i.e., the ratio of the number of adversarial blocks to the total number of blocks available in the canonical chain. Note that this ratio is not affected by the mining difficulty and is the same before and after applying a DAM. We normalize the two-week duration of an ideal epoch to 1 and the total reward for an epoch, $2016R$, to 1 as well. The duration of the starting epoch of the attack, namely epoch $e$, is denoted by $D_{e}$ which is greater than 1. The time-averaged profit of selfish mining in epoch $e$ can be obtained as follows:
(3) $\texttt{Profit}_{e}=\frac{\texttt{BR}^{\mathcal{A}}(\pi^{\texttt{selfish}})}{D_{e}}<\texttt{BR}^{\mathcal{A}}(\pi^{\texttt{selfish}})\enspace.$
Assume the adversary continues selfish mining in the upcoming epochs $E=\{e+1,e+2,\cdots\}$. Each epoch $e^{\prime}$ in set $E$ is preceded by an epoch that has gone under the selfish mining attack and thus has a reduced level of mining difficulty compared to epoch $e$. Since for any epoch $e^{\prime}\in E$ and its preceding epoch $e^{\prime}-1$, the effective mining power that contributes to the canonical chain is the same (both are under the selfish mining attack), the DAM at the end of epoch $e^{\prime}-1$ can properly adjust the mining difficulty for epoch $e^{\prime}$ so that the duration of epoch $e^{\prime}$ becomes 1. Therefore, the time-averaged profit of selfish mining in epoch $e^{\prime}\in E$ can be obtained as follows:
(4) $\texttt{Profit}_{e^{\prime}}=\texttt{BR}^{\mathcal{A}}(\pi^{\texttt{selfish}})\enspace.$
The equality above indicates that the time-averaged profit of selfish mining in epoch $e^{\prime}\in E$ equals the adversarial block ratio. However, since the time-averaged profit during the initial epoch of the attack is less than the adversarial block ratio, the total time-averaged profit from the start of the attack remains below the attacker’s block ratio. Over a long period of selfish mining, the time-averaged profit will eventually converge to the adversarial block ratio.
In the following, we analyze the minimum number of days required for a selfish mining attack to become profitable under the fixed block reward model, i.e., the first point at which the profit from selfish mining exceeds that of honest mining since the start of the attack. This period is referred to as the profit lag in (Grunspan and Pérez-Marco, 2023). Fig. 2 represents the minimum number of days required for a selfish mining attack to become profitable under the fixed block reward model. The figure is depicted for an adversary with different communication capabilities of $\gamma=0.5$, $0.75$, and $1$. To depict Fig. 2, we have assumed that the adversary follows the optimal selfish mining strategy introduced in (Sapirshtein et al., 2017), using which the adversary can achieve the highest possible block ratio. As can be seen in Fig. 2, the duration of the initial loss does not always decrease with an increase in the adversarial mining power, which seems counterintuitive. This is because the optimal selfish mining strategy introduced in (Sapirshtein et al., 2017) only aims to maximize the selfish miner’s block ratio in the canonical chain, indicating that this strategy optimizes the selfish miner’s profitability only after the difficulty adjustment mechanism is applied and the mining difficulty is reduced. However, as indicated in equation 3, the profitability of selfish mining in the first days of the attack depends not only on the block ratio achieved under a specific strategy but also on the duration of the initial epoch resulting from that strategy. Therefore, the optimal selfish mining strategy introduced in (Sapirshtein et al., 2017), which only considers optimizing the block ratio, is not necessarily the optimal strategy that an attacker can follow before the adjustment of the difficulty. For instance, consider a scenario where the adversarial communication capability is equal to 1, meaning the selfish miner can orphan one honest block for every adversarial block mined. In this situation, as the selfish miner’s mining power increases, more honest blocks will be orphaned during the initial epoch of selfish mining, resulting in a longer duration for this initial epoch. For detailed information regarding selfish mining profit lag, readers are referred to (Grunspan and Pérez-Marco, 2023) and (Sarenche et al., 2024b).
Figure 2. The minimum number of days required for a selfish mining attack to become profitable under the fixed reward model.
The main takeaway that Fig. 2 serves is that under the fixed block reward model, selfish mining takes a relatively long time (more than one ideal epoch or two weeks) to become profitable. Thus far, Bitcoin has not experienced a serious selfish mining attack, as the number of detected orphan blocks is negligible. These orphan blocks can likely be attributed to delays in publishing blocks in the network. One of the reasons that might make miners reluctant to engage in selfish mining is the relatively long initial period of the loss, which is caused by the infrequent application of the difficulty adjustment mechanism (once every two weeks). Although the infrequent DAM has some drawbacks, such as unstable block throughput during fluctuations in mining power, it can disincentivize mining power destruction attacks by prolonging the period of unprofitability.
3.3. Profitability of Pre-DAM Selfish Mining Under the Volatile Block Reward Model
A concerning point regarding Bitcoin’s security is that, due to the halving events, transaction fee rewards are expected to become comparable to or even exceed the protocol’s block rewards in the coming years, leading to more volatile block rewards. The volatility of the reward model is caused by the variation in transaction fees across different blocks. Under the volatile block reward model, the initial loss period of selfish mining can be shortened or even eliminated, regardless of the infrequent difficulty adjustment mechanism. Therefore, a selfish miner may see an immediate increase in its profit after starting the attack, suggesting that selfish mining could become a more significant threat to the Bitcoin network in the future.
While selfish mining under the fixed block reward model requires waiting for the difficulty adjustment to become profitable, selfish mining under the volatile block reward model can be profitable even before a DAM. Similar to the fixed block reward model, selfish mining does not increase the block generation rate before a DAM under the volatile block reward model. The key factor that makes the initial epoch of selfish mining profitable under the volatile model is that the block reward in this model is not fixed and can vary depending on the amount of fees paid by the transactions included in the block. Each block is typically filled with those transactions from the mempool that offer the highest transaction fees. The total transaction fees of a block depend on various factors [^2] Other variables such as transaction broadcasting rate, transaction fee rate, block space, etc. can also affect the total fee of a block., with block generation time playing a significant role in the profitability of selfish mining before a DAM. The block generation time is defined as the time gap between the proposal time of the last block in the canonical chain and the proposal time of the new block. There is competition among Bitcoin users seeking to have their transactions included in the blockchain quickly, requiring them to select an appropriate transaction fee. The longer it takes to generate a new block, the more intense the competition among transactions to be included in the next block becomes, resulting in a greater number of fee-valuable transactions being broadcast to the mempool. This implies that the total transaction fees included in a block increase as the block generation time grows. Let $\lambda_{1}$ [^3] Note that $\lambda_{1}=\lambda\alpha$, where $\lambda$ is the total block generation rate. and $R_{1}$ denote the adversarial canonical block mining rate and the average block reward under the volatile reward model when all the miners including the adversary mine honestly. Also, let $\lambda_{2}$ and $R_{2}$ denote the adversarial canonical block mining rate and the average block reward under the volatile reward model during the first epoch of selfish mining. As discussed in Section 3.1, the adversary’s time-averaged profit during the honest mining and the first epoch of selfish mining are equal to $\lambda_{1}R_{1}$ and $\lambda_{2}R_{2}$, respectively. By conducting a selfish mining attack, the adversarial canonical block generation rate decreases in the first epoch of the attack as some of the adversarial blocks may become orphaned, implying that $\lambda_{2}\leq\lambda_{1}$. However, unlike the fixed block reward model, $R_{2}$ is not necessarily equal to $R_{1}$. In the volatile model, as a result of the increase in the block generation time [^4] The network block generation time is inversely proportional to the total block generation rate. during the first epoch of the attack, the average block reward may increase, implying that $R_{2}\geq R_{1}$. The reason behind this increase in the block reward is intuitively illustrated in Fig. 3. Therefore, one of the time-averaged profits $\lambda_{i}R_{i}$ for $i\in\{1,2\}$ may be greater than the other depending on the adversary’s mining parameters and the mempool statistics. This shows that under the volatile reward block model, selfish mining can be profitable even in the first epoch of the attack.
Figure 3. Selfish mining under the volatile reward model. Assume that transaction fees can take values $i\in\{1,2,3,4\}$. The rate of transaction arrival for transactions with a fee value $i\in\{1,2,3,4\}$ is one transaction per 10 minutes. Furthermore, each block can include up to 4 transactions. Adversarial (honest) blocks are depicted in red (green). Under honest mining, the block generation time for all blocks is 10 minutes, and each adversarial block receives 10 units of fee. In the case of selfish mining, the block generation time for the second adversarial block increases. If the attack successfully orphans the middle honest block, the second adversarial block can collect an increased fee of 14 units.
3.4. Profitability of Selfish Mining After DAM Under the Volatile Block Reward Model
Let $e$ denote the initial epoch of selfish mining. As discussed in Section 3.2, after applying DAM at the end of epoch $e$, the mining difficulty for the subsequent epoch $e+1$ will be reduced. This reduction in difficulty implies that the adversary will be able to mine new blocks in a shorter time. Consequently, the average number of canonical adversarial blocks per unit of time will increase after the difficulty adjustment compared to before it. [^5] This is the case if the selfish miner has the minimum power share that makes selfish mining profitable. Regardless of whether transaction fees are comparable to the protocol reward, the increase in the average number of adversarial blocks per unit of time will raise the adversary’s reward per unit of time. Therefore, the adversary’s time-averaged profit increases after the difficulty adjustment in both the fixed and volatile reward models. An important question that arises in this context is which reward model—fixed or volatile—makes selfish mining more profitable after the difficulty adjustment. As demonstrated in the WeRLman paper (Bar-Zur et al., 2022), the minimum hash rate required for profitable deviation from honest mining after difficulty adjustment decreases in the volatile reward model. To explain how the volatile reward model lowers the deviation mining share threshold, we must focus on blocks with significantly higher rewards compared to those around them. These blocks may contain a whale transaction with an unusually large fee or may result from exceptionally long block generation times. Once such a block is mined, the adversary is incentivized to orphan the block, rather than extend it, in an attempt to steal the valuable transactions included within.
However, this reduction in the deviation mining share threshold does not necessarily imply that difficulty-adjusted selfish mining in the volatile reward setting is always more profitable for all miners and across all mempool patterns compared to the same attack under the fixed reward model. Before the difficulty adjustment, the main advantage of selfish mining in the volatile model compared to the fixed reward model is that it increases the average reward of an adversarial block after the attack is initiated. This occurs because selfish mining prolongs the canonical block generation time, allowing the next block to accumulate more valuable transaction fees. After the difficulty adjustment, however, the average block generation time in the canonical chain reverts to the standard 10-minute duration. This means that, after the difficulty adjustment, selfish mining under the volatile model loses its advantage of producing adversarial blocks with longer average generation times.
To compare difficulty-adjusted selfish mining profitability under the volatile and fixed reward models, we need to examine the ratio of the average canonical block reward before and after the attack in both models. Note that this ratio is 1 in the fixed reward model. However, in the volatile model, although the average block generation time remains the same before and after the attack, the different pattern of canonical block arrivals after the difficulty-adjusted selfish mining results in the average block reward not being necessarily the same before and after the attack. In the WeRLman paper (Bar-Zur et al., 2022), it is assumed that the rewards of all the normal blocks before and after the attack are the same; however, this assumption does not fully reflect what happens in practice. For instance, after the difficulty adjustment, the reduction in mining difficulty leads to shorter average block generation times for those canonical blocks that have no orphan blocks between them and their parent blocks, suggesting that these blocks tend to have lower fees compared to normal blocks generated before the attack.
To clarify that the average block reward under the volatile block reward model is not necessarily the same before and after difficulty-adjusted selfish mining, we first define the time-fee function and then present a theorem related to it. A time-fee function estimates the total transaction fees included in a block as a function of block generation time. Different mempool patterns can result in different time-fee functions.
Theorem 3.1.
If the time-fee function is strictly concave, the average fee of a canonical block under honest mining will be greater than the average fee of a canonical block under difficulty-adjusted selfish mining.
The proof of Theorem 3.1 is presented in Appendix C.
Observing Bitcoin mempool statistics (mem, 2024) reveals that there are frequent periods during which the corresponding time-fee function of the mempool is strictly concave, meaning the rate of increase in the block fee decreases over time. In other words, in most cases, a block $B_{1}$ with twice the generation time of block $B_{2}$ will accumulate less than twice the transaction fees of $B_{2}$. According to Theorem 3.1, under a concave time-fee function, the average canonical block fees after difficulty-adjusted selfish mining are lower than those before the attack. Therefore, since the ratio of the average block reward before and after the attack under the fixed reward model is 1, the percentage increase in profitability from difficulty-adjusted selfish mining under the fixed block reward can be higher than that under the volatile reward model. Note that this conclusion is based on the assumption of no or rare whale transactions. The presence of frequent whale transactions with substantial fees that do not follow the time-fee function pattern may offset the lower profitability of the attack in the volatile model resulting from the concavity of the time-fee function.
4. Simplified Volatile Reward Model
In this section, we provide a simplified approach to analyze mining strategies under the volatile reward model. We discuss how to implement a simplified environment that reflects the behavior of the Bitcoin mempool. Besides, we present a Markov Decision Process (MDP)-based tool that calculates a lower bound on the optimal time-averaged profit an adversary can achieve before a difficulty adjustment under the volatile reward model.
4.1. Simplified Mempool Implementation
In an MDP implementation, when the agent, which represents the adversary in our implementation, selects an action, the environment transitions to a new state and provides a corresponding reward. Under a fixed reward model, the reward is determined by the number of adversarial and honest blocks mined. However, in a volatile reward model, the block reward is dependent not only on the number of blocks but also on the transaction fees included in those blocks. To reflect the effect of transaction fees on the block rewards, our implemented environment encompasses an equation that calculates the transaction fee included in a block as a function of block generation time. We refer to this equation as the time-fee equation. Deriving a time-fee equation compatible with the Bitcoin environment is challenging, as transaction fees vary across different periods due to various reasons such as fluctuations in Bitcoin’s supply and demand dynamics. This variability indicates that the optimal mining strategy for an adversary can differ significantly across different periods. To maximize its profit, the adversary must accurately predict the time-fee equation for the targeted attack period.
For the simplified mempool implementation, we extracted historical data (Blo, 2024) on total transaction fees and block generation times for Bitcoin blocks mined across different periods. Using regression techniques, we derived the time-fee equation for the target period. This time-fee equation allows us to define the reward function for the MDP implementation. In Fig. 4, the total transaction fee of a block in BTC, denoted by $f(\cdot)$, is depicted as a function of its generation time in minutes, denoted by $t$, for two different periods. Each point represents a block, and the line represents the time-fee equation.
(a) October 8, 2023
(b) December 17, 2023
Figure 4. Linear time-fee equation
Fig. 4(a) corresponds to the Bitcoin blocks mined on October 8, 2023, which we denote as $\texttt{Day}_{1}$. The transaction fee per block during this day is relatively low. The corresponding linear regression line that represents the time-fee equation is $f_{1}(t)=0.0069t+0.0359$. As on average $1$ Bitcoin block is mined every $10$ minutes, the average transaction fee per block in this period is equal to $0.1049$ BTCs. Fig. 4(b) is related to the Bitcoin blocks mined on December 17, 2023, which we denote as $\texttt{Day}_{2}$. During this day, Bitcoin has experienced one of the periods with the highest transaction fee per block. The corresponding time-fee equation for this day is $f_{2}(t)=0.0659t+3.0457$, indicating that the average transaction fee per block is equal to $3.7052$ BTCs.
In this section, we modeled the Bitcoin mempool using linear regression, overlooking the concavity of the time-fee function, and employing only a single predictor of block generation time. To enhance mempool implementation, one can employ more advanced regression techniques as discussed in Appendix D.
4.2. MDP-Based Tool
In this section, we introduce an MDP-based tool to analyze the profitability of selfish mining before a difficulty adjustment under a highly simplified volatile block reward model. Our MDP-based tool is presented in (our, 2024). MDP can be used to determine the optimal policy when the number of possible states in the environment is limited to fewer than approximately $10^{6}$. Therefore, analyzing the volatile reward model with MDP requires simplifying the environment’s representation to fit within a limited number of states. Our MDP tool is designed to calculate a lower bound for selfish mining profitability under the volatile reward model, allowing us to determine the minimum amount of mining power required for a miner to be incentivized to deviate from the honest strategy.
In our simplified model, we consider two forks: an honest fork and an adversarial fork. The honest fork is publicly known to all miners, while the adversarial fork is known only to the selfish miner, who can strategically decide when to publish the blocks in the adversarial fork. To account for the volatile reward model, we classify the blocks in the adversarial fork into three categories:
  • Blocks with normal block generation time: These adversarial blocks have an average generation time of 10 minutes, as they are mined after another adversarial block, without any intervening honest blocks.
  • Blocks with extended block generation time: These adversarial blocks have longer generation times, as there is one honest block between them and their preceding adversarial blocks in the adversarial fork. On average, these blocks have a block generation time of 20 minutes.
  • Blocks with highly extended block generation time: These adversarial blocks have even longer generation times, with at least 2 honest blocks intervening between them and their preceding adversarial blocks in the adversarial fork. On average, these blocks have a block generation time of 30 minutes or more.
Blocks with extended or highly extended generation times offer higher fee rewards compared to those with normal times. Thus, to achieve the optimal mining strategy in this simplified model, the adversary should aim to maximize the number of extended and highly extended blocks. The reason we consider only up to blocks with a 30-minute generation time is to limit the state space.
The actions available to the adversary are similar to those introduced in (Sapirshtein et al., 2017) and include: override, adopt, match, and wait. In our simplified environment, each state is represented by five elements: $(l_{\mathcal{A}},l_{\mathcal{H}},G,\texttt{Latest},\texttt{Match})$. Here, $l_{\mathcal{A}}$ and $l_{\mathcal{H}}$ denote the lengths of the adversarial and honest forks, respectively, and can range from 0 to maxForkLen, where maxForkLen represents the maximum fork length in this environment. If the adversarial fork reaches maxForkLen, the selfish miner should publish the fork; if the honest fork reaches this length, the adversary should adopt the honest fork. The parameter maxForkLen limits the state space to a finite number of states. $G$ is an array of length maxForkLen that stores information about the block generation times in the adversarial fork. Each element in $G$ can be 0, 1, or 2: if the $i^{\text{th}}$ element is 0, the corresponding block in the adversarial fork has a normal generation time; if the $i^{\text{th}}$ element is 1, the block has an extended generation time; and if the $i^{\text{th}}$ element is 2, the block has a highly extended generation time. Latest is an element indicating how many honest blocks have been mined since the last adversarial block. It can take values 0, 1, or 2, denoting that the next adversarial block has a normal, extended, or highly extended generation time, respectively. Lastly, Match is a boolean element indicating whether the match action is currently in progress.
4.3. Results and Implications: Selfish Mining Under a Simplified Volatile Model
To analyze selfish mining profitability over different periods, we first define base fee and fee increase rate. The base fee is the intercept of a linear time-fee equation that represents the average amount of remaining transaction fees available in the transaction pool immediately after a block is mined. The fee increase rate is the slope of a linear time-fee equation that represents the increase in the collected transaction fee of a block per unit of time. At $t=0$, the transaction reward of the next potential block is equal to the base fee. This reflects reality, as there are always some low-value transactions in the mempool that miners can use to fill their blocks. As block generation time increases, the collected transaction fee rises by the fee increase rate for each additional unit of time due to the arrival of new transactions.
The impact of the base fee on selfish mining profitability is similar to that of the block protocol reward, as a block always receives the base fee in addition to the protocol reward, regardless of the block generation time. The main factor that results in selfish mining profitability before difficulty adjustment under the volatile reward model is the fee increase rate. The higher the fee increase rate, the more profitable selfish mining would be. The intuition, as discussed in Section 3.3, is that selfish mining increases the block generation time before a difficulty adjustment mechanism. This increase in the block generation time can raise the collected fee of the adversarial block proportional to the fee increase rate. To intuitively analyze the effect of the time-fee equation on selfish mining profitability before the difficulty adjustment, we define the term orphaning fee ratio. This term is defined as the ratio of the average reward received by a block with a generation time of 20 minutes to the average reward received by a block with a generation time of 10 minutes. The former block represents an adversarial block that has managed to orphan its previous block, while the latter represents a block mined following the honest strategy. The greater the orphaning fee ratio exceeds 1, the more profitable selfish mining becomes. If the orphaning fee ratio is close to 1, the block reward does not increase substantially over time, implying that selfish mining may not be significantly profitable.
We compare selfish mining profits before difficulty adjustment during $\texttt{Day}_{1}$ and $\texttt{Day}_{2}$ introduced in Section 4.1, under two conditions: when the protocol reward is still substantial and after it has converged to zero. We normalize the time-averaged profit of the entire network to 1, indicating that if all miners mine honestly, the time-averaged profit for each miner will equal their mining share. In Fig. 5, the lower bound of the time-averaged profit that a selfish miner can achieve before the difficulty adjustment on $\texttt{Day}_{1}$ and $\texttt{Day}_{2}$, under both protocol rewards of 0 and 3.125 BTC, is depicted as a function of selfish miner’s mining share. This graph corresponds to the selfish miner with a communication capability equal to $1$. We proceed by analyzing the results depicted in Fig. 5. Let $R(t)$ denote the total reward of a block as a function of block generation time and $r^{\texttt{orphan}}$ denote the orphaning fee ratio. Assuming a protocol reward of $3.125$ BTCs per block, the block reward-time equation during $\texttt{Day}_{1}$ and $\texttt{Day}_{2}$ can be obtained as $R_{1}(t)=0.0069t+3.1609$ and $R_{2}(t)=0.0659t+6.1707$, respectively. In $\texttt{Day}_{1}$, where the transaction fee is negligible compared to the protocol reward, the orphaning fee ratio is $r_{1}^{\texttt{orphan}}=1.0213$, making selfish mining not substantially profitable before the difficulty adjustment. However, on $\texttt{Day}_{2}$, where the transaction fee is comparable to the block reward, the orphaning fee ratio $r_{2}^{\texttt{orphan}}=1.0965$. As $r_{2}^{\texttt{orphan}}>r_{1}^{\texttt{orphan}}$, under the assumption of $3.125$ BTCs protocol reward per block, selfish mining is more profitable before the difficulty adjustment on $\texttt{Day}_{2}$ compared to $\texttt{Day}_{1}$.
Figure 5. A lower bound for the time-averaged profit before difficulty adjustment.
Considering the scenario that the protocol reward has converged to 00, the block reward-time equations would be the same as the time-fee equations. Under this scenario, the orphaning fee ratios are $r_{1}^{\texttt{orphan}}=1.6578$ for $\texttt{Day}_{1}$ and $r_{2}^{\texttt{orphan}}=1.1779$ for $\texttt{Day}_{2}$, respectively. Although the transaction fees on $\texttt{Day}_{1}$ are much lower than transaction fees on $\texttt{Day}_{2}$, as $\texttt{Day}_{1}$ has a higher orphaning fee ratio, miners are more incentivized to conduct selfish mining on $\texttt{Day}_{1}$ once the protocol reward converges to 00.
These observations conclude that selfish mining before a difficulty adjustment becomes profitable on days when the ratio of the fee increase rate to the base reward (base fee plus protocol reward) is substantial. To meet this condition with a significant protocol reward, transaction fees must also be high and comparable to the protocol reward. However, if the protocol reward becomes negligible, the Bitcoin network becomes more vulnerable to selfish mining on days when the rate of increase in transaction fees is high, even if those days do not coincide with high transaction fees per block.
5. Enhanced Volatile Reward Model
This section introduces our Asynchronous Advantage Actor Critic (A3C)-based implementation designed to derive a near-optimal mining policy within a more realistic volatile reward model. A brief discussion of the A3C algorithm is provided in Appendix B. Our implementation codes are presented in (our, 2024). The environment used by the A3C-based tool includes a transaction mempool, allowing us to realistically simulate the behavior of the Bitcoin mempool. Our A3C-based implementation can analyze mining attacks such as selfish mining and undercutting, both before and after the difficulty adjustment mechanism.
As discussed in Appendix B, A3C is a reinforcement learning algorithm where the agent (the payoff-maximizing miner) interacts with the environment (the Bitcoin blockchain). Through these interactions, the agent transitions between different states, each containing the necessary information to represent both blockchain forks and mempool conditions. At each state, the adversary can choose from various possible actions, with each action offering different rewards. A detailed discussion of the possible action set and state representation in our implementation is presented in Appendix E.1.
5.1. Enhanced Mempool Implementation
In the simplified model, we estimate the time-fee equation using only data on block transaction fees and generation times. However, the way transaction fees impact block rewards is more complex in practice. To better estimate how transaction fees fluctuate over time, we can improve the model by incorporating the behavior of the Bitcoin mempool. The mempool is a storage space where unconfirmed transactions are stored before they are included in a block. To increase the chance of inclusion in a block, each Bitcoin transaction pays a fee to the block miner. The total fee that a transaction pays is calculated based on the transaction’s weight, which is measured in virtual bytes (vBytes). Virtual bytes are a weight unit introduced by the SegWit upgrade. The fee is typically expressed in satoshis per virtual byte (sat/vByte), where each satoshi is $10^{-8}$ BTC. Each Bitcoin block has a size limit of 1 virtual megabytes (vMB) (Bit, 2024). Miners prioritize filling the block with the most valuable transactions from the mempool, where the term valuable refers to transactions that offer the highest fees in satoshis per virtual byte. In the context of the enhanced volatile reward model, we define the base fee for a given mempool period as the fee for which there is at least 1 virtual megabyte of transactions offering that fee or higher at all times during that period.
Our implementation categorizes the transactions available in the mempool into $N^{\text{range}}$ sat/vByte ranges, denoted as
$\{\texttt{fee}_{0},\texttt{fee}_{1},\ldots,\texttt{fee}_{N^{\text{range}}-1}\}\enspace,$
where the fee levels are ordered in ascending sequence. A transaction paying $f$ sat/vByte, where $\texttt{fee}_{j}\leq f<\texttt{fee}_{j+1}$, is assigned to the range corresponding to $\texttt{fee}_{j}$. $\texttt{fee}_{0}$ denotes the base fee, implying that all transactions available in the mempool pay at least $\texttt{fee}_{0}$ sat/vByte. In our implementation, we assume an unlimited number of transactions paying the base fee. The weight of transactions in the range $\texttt{fee}_{j}$ increases over time according to a specific equation, called the weight-time equation, which is estimated based on the transaction weight growth in the range $\texttt{fee}_{j}$. Whenever a new block is mined, a total weight equivalent to 1 virtual megabyte is deducted from mempool, starting from the highest sat/vByte ranges. To estimate the weight increase function for each sat/vByte range over a given time period, we analyzed mempool statistics during that period using the information presented in (mem, 2024). Using these statistics, we applied regression techniques to derive the weight-time equation for each sat/vByte range.
The estimated weight-time curves based on mempool statistics from December 17, 2023, are shown in Fig. 6. To depict Fig. 6, the transactions in mempool are categorized into five ranges of sat/vByte, with the base fee equal to $200$ sat/vByte. This base fee is chosen because, at all times on December 17, 2023, there was at least 1 virtual MB of transactions in the mempool paying a fee higher than $200$ sat/vByte. This indicates that, regardless of the block generation time, a block mined on this day could collect at least 200 million satoshis (2 BTC). An interesting observation from Fig. 6 is that, although the weights of transactions in all sat/vByte ranges are increasing over time, the rate of increase for weights in the lower sat/vByte ranges decreases over time (indicating a concave weight-time equation). In contrast, the rate of increase for weights in the higher sat/vByte ranges accelerates (indicating a convex weight-time equation). This observation aligns with the competition that exists in practice among transactions for inclusion in blocks. As the generation time for the next potential block increases, this competition intensifies, where newer transactions need to submit higher fees to maintain their chances of being included in the next block.
Figure 6. Total transaction weight for different ranges of sat/vByte as a function of block generation time (December 17, 2023).
5.2. Objective Function
The training process aims to enable agents to identify which action maximizes the objective function at a given state $s$. The objective function is defined in terms of the parameters that the agent, the adversary in our case, seeks to maximize. In our implementation, this objective is to maximize the mining time-averaged profit. We discuss how the objective function can be defined both before and after the difficulty adjustment. In our implementation, at each time step $t$, a block is generated either by the adversary or by the honest miners. However, depending on the adversarial strategy, the block generated at time step $t$ may not immediately be added to the canonical chain. This block could be added to the canonical chain or even be identified as orphaned at a future time step. Let $N_{\mathcal{A}}(t;\pi)$, $N_{\mathcal{H}}(t;\pi)$, and $R_{\mathcal{A}}(t;\pi)$ denote the number of adversarial blocks, the number of honest blocks, and the total reward included in the adversarial blocks, respectively, added to the canonical chain at time step $t$ under policy $\pi$. Also, let $N_{\mathcal{T}}(t;\pi)$ denote the total number of blocks both added to and removed (orphaned) from the canonical chain at time step $t$ under policy $\pi$. For each time step $t$, we have $N_{\mathcal{T}}(t;\pi)\geq N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)$. After the state transition at time step $t$, the environment returns the next state along with the following information to the agent: $\big{(}R_{\mathcal{A}}(t;\pi),N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi),N_{\mathcal{T}}(t;\pi)\big{)}$.
We denote by $\mathbf{t}^{B}$ the average block generation time. We normalize the average block generation time achieved while all the miners including the adversary mine honestly to $\mathbf{t}^{B}_{\texttt{ideal}}=1$. Let us ignore the mining cost since, as long as the adversary does not turn off some of its miners, its mining cost is not affected by the chosen strategy. Therefore, the time-averaged profit can be defined as follows:
(5) $\texttt{Profit}^{\mathcal{A}}(\pi)=\lim_{T\to\infty}{\frac{\sum_{t=0}^{T}{R_{\mathcal{A}}(t;\pi)}}{\mathbf{t}^{B}\cdot T}}\enspace.$
Depending on the mining difficulty and the adversary’s strategy, the average block generation time $\mathbf{t}^{B}$ may vary, leading to different objective functions before and after the difficulty adjustment.
5.2.1. Before the difficulty adjustment
As discussed in Section 3.1, before a difficulty adjustment, the adversary’s mining strategy does not affect the average block generation time, and thus, $\mathbf{t}^{B}=1$. Note that while the canonical block generation time may differ, the average duration of each time step in our environment before a difficulty adjustment is determined by the average of both canonical and non-canonical block generation times, which remains unchanged. Therefore, the objective function before the difficulty adjustment can be defined as follows:
(6) $O^{\mathcal{A}}(\pi)=\lim_{T\to\infty}{\frac{\sum_{t=0}^{T}{R_{\mathcal{A}}(t;\pi)}}{T}}\enspace.$
To find the policy that maximizes the optimization function above, the reward that agents receive after a transition at time step $t$, referred to as the time step reward and denoted by $r_{t}$, is set to $R_{\mathcal{A}}(t;\pi)$. During our implementation, it was found that including the mining cost in the reward function (as a constant negative reward added after each time step to account for the passage of time) and normalizing the rewards helps the agent converge faster and in a more stable manner. Our implementation uses the following step reward function:
(7) $r_{t}=\frac{R_{\mathcal{A}}(t;\pi)}{R^{\texttt{norm}}}-\texttt{cost}\enspace,$
where $R^{\text{norm}}$ denotes the total distributed reward on average per $\mathbf{t}^{B}$ units of time if all miners mine honestly, and cost denotes the constant cost per time step.
5.2.2. After the difficulty adjustment
After the difficulty adjustment, the mining difficulty will be modified, resulting in a decrease in the block generation time. The average block generation time after difficulty adjustment and up to including time step $T$ can be obtained as follows:
(8) $\begin{split}&\mathbf{t}^{B}=\frac{\sum_{t=0}^{T}{N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)}}{\sum_{t=0}^{T}{N_{\mathcal{T}}(t;\pi)}}\cdot\mathbf{t}^{B}_{\texttt{ideal}}=\\ &\frac{\sum_{t=0}^{T}{N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)}}{T}\enspace.\end{split}$
The second equality in the equation above is achieved as $\mathbf{t}^{B}_{\texttt{ideal}}=1$ and $\sum_{t=0}^{T}{N_{\mathcal{T}}(t)}=T$. Based on the value of $\mathbf{t}^{B}$ and the time-averaged profit formula outlined in Equation 5, we define the objective function after difficulty adjustment as follows:
(9) $O^{\mathcal{A}}(\pi)=\lim_{T\to\infty}{\frac{\sum_{t=0}^{T}{R_{\mathcal{A}}(t;\pi)}}{\sum_{t=0}^{T}{N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)}}}\enspace.$
Since the objective function defined in equation 9 is non-linear, we cannot naively assign the time step reward as is done in Section 5.2.1. To address this challenge, we adopt the approach introduced in (Sapirshtein et al., 2017), which converts the non-linear optimization function into a linear one. We set the time step reward as follows:
(10) $r_{t}(\rho)=R_{\mathcal{A}}(t;\pi)-\rho\big{(}N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)\big{)}\enspace,$
where $\rho$ represents the time-averaged profit the adversary earns by following the optimal strategy, i.e., $\rho=O^{\mathcal{A}}(\pi^{*})$, where $\pi^{*}$ denotes the optimal strategy. At the start of training, the actual value of $\rho$ is unknown. Therefore, the training agents initialize with $\rho=O^{\mathcal{A}}(\pi^{\mathcal{H}})$, where $\pi^{\mathcal{H}}$ denotes the honest strategy. Certain test agents are tasked with measuring the time-averaged profit under the policy trained by the training agents at specific time steps during the training process. If this profit exceeds the highest time-averaged profit achieved so far, they prompt the training agents to update $\rho$ to the new highest time-averaged profit. For the detailed analysis of the non-linear objective function conversion readers are referred to (Sapirshtein et al., 2017). We briefly provide the intuition behind setting the time step reward as defined in equation 10 in Appendix F.1.
Another critical point is that the average time step duration after difficulty adjustment needs recalibration in the environment. For details on how our implementation measures the time step duration, readers are referred to Appendix F.2.
5.3. Results and Implications: Mining Attacks Under an Enhanced Volatile Model
This section presents results from our A3C-based implementation (our, 2024), analyzing mining attacks under the mempool patterns observed during different periods. Information regarding our A3C-based tool implementation is provided in E.4. To compare the time-averaged profits of a mining strategy $\pi$, we use the concept of the percentage increase in the miner’s time-averaged profit, defined as follows:
(11) $\texttt{PI}^{\mathcal{A}}(\pi)=\frac{\texttt{Profit}^{\mathcal{A}}(\pi)-\texttt{Profit}^{\mathcal{A}}(\pi^{\mathcal{H}})}{\texttt{Profit}^{\mathcal{A}}(\pi^{\mathcal{H}})}*100\%\enspace,$
where $\texttt{Profit}^{\mathcal{A}}(\pi^{\mathcal{H}})$ represents the time-averaged profit achieved by the adversary if all miners, including the adversary, follow the honest strategy.
In this section, we analyze the profitability of selfish mining before the difficulty adjustment, as well as the single-block undercutting attack. The results related to the profitability of selfish mining after the difficulty adjustment is presented in Appendix G.
5.3.1. Mining attacks before a difficulty adjustment.
In Figure 7, the percentage increase in the adversary’s time-averaged profit before a difficulty adjustment is shown as a function of the adversary’s mining share. This figure is depicted based on the following assumptions: i) the protocol reward is set to $3.125$ BTC, ii) the mempool pattern observed on December 18, 2023, between 23:00 and 24:00 CET, has occurred, iii) the adversary’s communication capability equals 1.
Figure 7. The percentage increase in time-averaged profit (December 18, 2023; protocol reward: $3.125$ BTC).
Figure 7 presents the results obtained from both our MDP implementation and A3C implementation. As shown in Figure 7, the strategy derived from the A3C implementation can outperform the strategy obtained from the MDP. Furthermore, as the adversary’s mining share increases, the gap between the results obtained from these two approaches grows. As can be seen in Figure 7, even with a protocol reward of $3.125$ BTC, on a day where the average transaction fee per block is comparable to the protocol reward (the average transaction fee per block on December 18, 2023, was $2.32$ BTC (exp, 2024)), the selfish mining attack can lead to profitability immediately after the attack begins.
Similarly, Figure 8 illustrates the percentage increase in the adversary’s time-averaged profit before a difficulty adjustment, using the same assumptions as in Figure 7, with the exception that the protocol reward is set to zero. By comparing Figures 8 and 7, we observe that as the protocol reward decreases, deviations from the honest strategy become increasingly profitable, posing a potential threat to Bitcoin’s progress in the near future.
Figure 8. The percentage increase in time-averaged profit (December 18, 2023; protocol reward: 00 BTC).
Although Figure 8 shows that in the transaction-fee era, the same mempool pattern as occurred on December 18, 2023, can be concerning, there are some other days where the threat of adversarial attack is more alarming. Figure 9 represents the percentage increase in profit before the difficulty adjustment from applying both selfish mining and undercutting attacks simultaneously under the mempool patterns of two different periods: December 18, 2023, between 23:00 and 24:00 CET (the same period as in Figures 8 and 7), and October 15, 2023, between 8:00 and 9:00 CET. The figure is depicted as a function of different ratios of petty-compliant miners under the following assumptions: i) the protocol reward equals zero, ii) the adversarial mining share is $\frac{1}{3}$, and iii) the adversary has a normal communication capability of 50%. As can be seen in Figure 9, even if all the honest miners are altruistic, the deviation still yields a substantial immediate profit for the adversary under the mempool pattern of October 15, 2023. By increasing the ratio of petty-compliant miners in the system, the deviation can become even more profitable. The main reason for the higher profitability of mining attacks under the mempool pattern of October 15, 2023, compared to that of December 18, 2023, is the low level of base fee available in the mempool. Between 8:00 and 9:00 CET on October 15, 2023, the base fee was 1 sat/vByte (mem, 2024), implying that during this period, miners needed to fill at least some part of their blocks with transactions offering only 1 sat/vByte. In contrast, the base fee was 80 sat/vByte on December 18, 2023, between 23:00 and 24:00 CET (mem, 2024). This comparison shows that a higher base fee level for transactions available in the mempool can help mitigate mining attacks in the transaction-fee era.
Figure 9. The percentage increase in time-averaged profit (protocol reward: 00 BTC).
5.3.2. Single block undercutting attack.
In our A3C-based implementation, we treat the undercut action as a continuous action, where the agent must not only decide whether to take this action but also determine the duration for which it should perform undercutting. Once the designated time has passed, the agent should stop undercutting the tip block and begin mining on top of it, as a sufficient number of transactions will have arrived in the system. For more information on the continuity of the undercut action, readers are referred to Appendix E.2.
To measure the profitability of the single-block undercutting attack, we have limited the adversary’s possible action set, as outlined in Appendix E.3. Figure 10 shows the percentage increase in time-averaged profit achieved by performing the undercutting attack. The figure is depicted based on mempool patterns observed over five different days between 8:00 and 9:00 CET, under the following assumptions: i) the protocol reward is set to 0, and ii) the adversary’s communication capability equals 1.
Figure 10. The profit percentage increase of undercutting.
The base fees for the corresponding periods are also depicted in Figure 10. Similar to selfish mining, a mempool with lower base fees is more vulnerable to the undercutting attack. However, other factors such as the transaction arrival rate also affect mining attack profitability. For example, comparing the mempool statistics from August 28, 2024, and September 6, 2024, we observe that although the base fee levels are nearly the same, the average weight of transactions arriving per 10 minutes during the former period is $0.98$ virtual megabytes, while in the latter period it is $0.65$ virtual megabytes (mem, 2024). As fewer transactions were arriving in the mempool on September 6, 2024, it becomes more profitable to undercut and steal transaction fees from another block during the same mempool pattern as this period.
6. Conclusion
In this paper, we analyzed the impact of volatility in block rewards on mining strategies in Bitcoin. Under the fixed block reward model, which is primarily suitable for modeling Bitcoin’s behavior during its first decade of existence, mining attacks such as selfish mining need to continue for at least one difficulty epoch (approximately two weeks) to become profitable. However, using historical Bitcoin statistics, we have demonstrated that under a volatile block reward model, when transaction fees become a major source of incentive, adversarial strategies can be profitable immediately, bypassing the initial loss period. The threat of deviant strategy profitability is concerning as the protocol reward diminishes to zero, and even under the current protocol reward, on days when average fees per block are comparable to the protocol reward.
Once assigning a transaction fee for their transactions, Bitcoin users should view their paid fee not only as a reward to compensate miners but also as a factor that can significantly impact the security of their valued blockchain. To ensure the secure progress of Bitcoin, users should provide a stable source of incentives for miners, so they have no, or at least minimal, motivation to deviate from the Bitcoin protocol.
Appendix A Mining attacks
A.1. Selfish mining
In the selfish mining attack, introduced by Eyal et al. in (Eyal and Sirer, 2018), a selfish miner withholds a newly mined block $B_{1}$ instead of immediately publishing it to the network. By keeping $B_{1}$ secret, the selfish miner creates a private fork one block ahead of the public chain. While other miners continue mining on the public chain, the selfish miners work to extend their private forks. If a selfish miner mines a second block $B_{2}$ on top of its private fork, its chain gains a two-block lead over the public chain. When an honest miner mines a new block $B_{3}$ on top of the public chain, the selfish miner can reveal its secret fork. Since the selfish miner’s chain, extended by blocks $B_{1}$ and $B_{2}$, is longer compared to the public chain, extended by block $B_{3}$, the selfish miner’s chain becomes the canonical chain. Consequently, block $B_{3}$, mined by honest miners, becomes orphaned and is excluded from the canonical chain. However, it is not necessarily the case that the selfish miner can always achieve a two-block lead. If the selfish miner keeps its block secret and honest miners mine the next block, a fork race occurs between two forks of the same height, with only one fork becoming the canonical chain. The probability of winning this race depends on how quickly the selfish miner can propagate its blocks through the network, encouraging other miners to build on top of the selfish miner’s fork. This attack can increase the selfish miner’s block reward relative to other miners. In (Sapirshtein et al., 2017), the authors implemented an MDP-based tool to determine the optimal selfish mining strategy that a selfish miner can follow in a setting where the reward for all blocks is the same, referred to as the fixed block reward model. In (Grunspan and Pérez-Marco, 2018), the authors argued that the block ratio is not a good benchmark for analyzing selfish mining profitability and that one should use the time-averaged profit definition for this analysis. They demonstrated that selfish mining during the first difficulty epoch cannot be more profitable than honest mining, resulting in an initial loss period for selfish mining. However, from the subsequent epoch onward, the attack becomes profitable as the difficulty decreases due to the selfish mining attack. To reduce the initial loss period, the authors in (Negy et al., 2020) introduced intermittent selfish mining, where the selfish miner alternates between one epoch of selfish mining and one epoch of honest mining. The authors in (Grunspan and Pérez-Marco, 2023) introduced the term profit lag to describe the point at which the attack consistently becomes more profitable than honest mining in order to analyze the initial period of loss. They argued that although intermittent selfish mining’s profitability may exceed that of honest mining at certain points in the early cycles of the attack, its profit lag is actually longer than the profit lag of standard selfish mining. In (Sarenche et al., 2024b), the authors introduced smart intermittent selfish mining, where the selfish miner performs selfish mining for half of each block. Smart intermittent selfish mining can outperform the intermittent selfish mining introduced in (Negy et al., 2020) and reduces the profit lag. The selfish mining attack is not specific to the Bitcoin network and can be applied to any longest-chain-based blockchain. As analyzed in (Brown-Cohen et al., 2019; Sarenche et al., 2024a), selfish mining can be even more destructive in the context of longest-chain Proof-of-Stake protocols due to factors such as proposer predictability and the nothing-at-stake phenomenon.
A.2. Undercutting
The undercutting attack, introduced by Carlsten et al. in (Carlsten et al., 2016), can be profitable in scenarios where a significant portion of the mining power is petty-compliant. Petty-compliant miners are those who may deviate from the honest mining strategy to earn higher profits. As the protocol reward approaches zero, the undercutting attack poses an increasingly serious threat to the stability of the Bitcoin network. In this attack, when an undercutter becomes aware of a new block $B_{1}$ that extends the longest chain, it may choose not to accept it. If $B_{1}$ contains the majority of the transaction fees available in the mempool, the undercutter might be disincentivized from mining a relatively empty block on top of it. Instead, the undercutter may opt to mine on the parent of $B_{1}$ to create a competing block. The key aspect of the undercutting attack is that the undercutter should include fewer transaction fees in the competing block compared to $B_{1}$, leaving a substantial portion of the fees in the mempool for the future block. This strategy of generously leaving some transactions is crucial for the attack’s success because if the undercutter successfully mines the competing block $B_{2}$, a fork will occur in the network between blocks $B_{1}$ and $B_{2}$. In the event of such a fork, petty-compliant miners would prefer to mine on top of the block that offers the most available transaction fees, rather than the oldest block—–in this case, the competing block $B_{2}$. Attracting the mining power of rational miners to $B_{2}$ increases the undercutter’s chances of winning the fork. If the next block is mined on top of the competing block $B_{2}$, the undercutter has successfully undercut block $B_{1}$. Carlsten et al. find that undercutting can become the equilibrium strategy for miners, leading to instability as miners undercut each other. However, this result is based on a model disregarding the block size limit. In paper (Gong et al., 2022), the authors have analyzed the undercutting attack with consideration of the block size limit and derived closed-form conditions on the pending transaction set in the mempool that make undercutting profitable. They have also proposed an alternative transaction selection rule to counter undercutting; instead of fitting all available transactions into a block, miners include only a portion of the pending transaction fees to mitigate the risk of being undercut.
Appendix B Asynchronous Advantage Actor-Critic (A3C)
Reinforcement learning (RL), a branch of artificial intelligence (AI), aims to develop autonomous agents that learn optimal policies to maximize long-term rewards through iterative interactions with their environment (Alzubaidi et al., 2021). Asynchronous Advantage Actor-Critic (A3C) (Mnih et al., 2016) is a policy gradient method within the RL domain that simultaneously optimizes both the policy and value function estimation. The policy, denoted by $\pi$, guides the agent’s actions in the environment by assigning probabilities to each possible action in a given state, indicating the likelihood of the agent choosing that particular action. The value function at a given state $s$, denoted by $V(s;\pi)$, estimates the expected cumulative reward that an agent can obtain starting from state $s$ and following a particular policy $\pi$ thereafter, formally defined as follows:
(12) $V(s;\pi)=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}\mid s_{t}=s,\pi\right]\enspace,$
where $s_{t}$ is the state visited at step $t$, $r_{t}$ is the reward agent receives at step $t$, and $\gamma$ is the discount factor. A3C uses multiple agents that interact with their own instances of the environment in parallel. These agents run independently and asynchronously to explore different parts of the state space simultaneously, reducing the correlation between experiences and improving exploration. Unlike many reinforcement learning algorithms, such as deep Q-learning (DQN), which rely on experience replay (storing and reusing past experiences), A3C updates the network in real-time as experiences are generated. The diversity and independence of experiences gathered by multiple agents compensate for the lack of experience replay. A3C uses 1-to-$n$-step returns to update both the policy and the value function. These updates occur after every $T$ steps or when a terminal state is reached. The value loss for the trajectory over the time steps $t\in\{0,1,\cdots,T-1\}$, denoted by $L_{v}$, is defined as the Mean Squared Error (MSE) between the value function and the expected return as follows:
(13) $L_{v}(\theta_{v})=\frac{1}{2}\mathbb{E}_{t}\Big{[}\big{(}R_{t}-V(s_{t};\theta_{v}))^{2}\Big{]}\enspace,$
where $R_{t}$ is the expected return in step $t$ with respect to the updating time step $T$ and is defined as follows:
(14) $R_{t}=\sum_{k=0}^{T-1-t}{\gamma^{k}r_{t+k}}+\gamma^{T-t}V(s_{T};\theta_{v})\enspace.$
The main goal of the policy update in A3C is to improve the policy $\pi$ such that actions with higher advantages are more likely to be chosen. The policy loss function for the trajectory over the time steps $t\in\{0,1,\cdots,T-1\}$, denoted by $L_{\pi}$ is formulated as follows:
(15) $L_{\pi}(\theta_{\pi})=\mathbb{E}_{t}\Big{[}\log\pi(a_{t}|s_{t})\cdot A_{t}\Big{]}\enspace,$
where $\pi(a_{t}|s_{t})$ represents the probability of taking action $a_{t}$ at state $s_{t}$ according to the current policy, and $A_{t}$ is the advantage function defined as $R_{t}-V(s_{t})$. Note that while the policy parameters ($\theta_{\pi}$) and value function parameters ($\theta_{v}$) are typically treated as separate for generality, in practice, some of these parameters are often shared.
Appendix C Proof of Theorem 3.1
Proof.
Let $f(t)$ denote the fee of a block with a block generation time of $t$, where $f(t)$ is strictly concave. Let $\lambda$ and $\lambda^{\prime}$ denote the total block generation rates under honest mining and difficulty-adjusted selfish mining, respectively. Note that the block generation time follows an exponential distribution. The canonical block generation rate under honest mining is also equal to $\lambda$. Therefore, the average canonical block fee under honest mining, denoted by $r_{\mathcal{H}}$, can be obtained as follows:
[…]
We divide canonical blocks mined under difficulty-adjusted selfish mining into groups $\{G_{1},G_{2},G_{3},\cdots\}$. A canonical block $B$ belongs to group $G_{i}$ for $i\in\{1,2,3,\cdots\}$ if the number of orphaned blocks mined between the parent of block $B$ and block $B$ itself is $i-1$. The average canonical block generation rate of blocks belonging to group $G_{i}$ is equal to $\frac{\lambda^{\prime}}{i}$. Let $a_{i}$ denote the ratio of the number of blocks belonging to group $G_{i}$ to the total number of canonical blocks, where $\sum_{i=1}^{\infty}{a_{i}}=1$. As the average canonical block generation rate under difficulty-adjusted selfish mining is also equal to $\lambda$, we have:
[…]
By calculating the result of integral, we have:
(18) $\frac{1}{\lambda}=\frac{\sum_{i=1}^{\infty}{a_{i}i}}{\lambda^{\prime}}\Longrightarrow\frac{\lambda^{\prime}}{\lambda}=\sum_{i=1}^{\infty}{a_{i}i}\enspace.$
The average fee of a canonical block under difficulty-adjusted selfish mining, denoted by $r_{\mathcal{S}}$, can be obtained as follows:
[…]
By making a variable substitution of form $\lambda t=\frac{\lambda^{\prime}}{i}t^{\prime}$, we obtain:
[…]
As function $f(t)$ is strictly concave, we have:
[…]
Using the result obtained in equation 18, we can write the inequality above as follows:
[…]
The right-hand side in the inequality above is equal to $r_{\mathcal{H}}$. Therefore, we obtain $r_{\mathcal{S}}<r_{\mathcal{H}}$. ∎
Appendix D Advanced Regression Techniques
In Fig. 4 presented in Section 4.1, a linear regression model is used to represent the time-fee equation. However, the block transaction fee does not necessarily increase linearly with the block generation time. This occurs when the rate of fee increase is not constant and decreases as the block generation time increases. In such cases, using alternative regression techniques rather than a linear model can provide a time-fee equation that better fits the behavior observed in the Bitcoin mempool. In Fig. 11(a), we compare linear regression with a curve regression model given by $f(t)=0.6414\;\texttt{ln}(t+6.5209)-0.8419$ to derive the time-fee equation for Bitcoin blocks mined on May 3, 2023. The coefficient of determination (R-squared) for the curve regression is $0.6243$, while for the linear regression it is $0.5180$. This indicates that the curve regression provides a better fit compared to the linear regression, suggesting that the fee increase rate diminishes with longer block generation times.
Another technique that can lead to a more realistic mempool implementation is using multivariate regression instead of univariate regression. In univariate regression, the block transaction fee is estimated based solely on a single predictor, namely the block generation time. In contrast, multivariate regression allows us to account for additional predictors, such as the generation time of the parent block, in determining the block transaction fee. In Fig. 11(b), we depict the regression plane for the block transaction fee on March 18, 2024, based on two predictors: the block generation time and the parent block generation time. The resulting time-fee equation is $f(t_{1},t_{2})=0.1697+0.0079*t_{1}+0.0046*t_{2}$, where $t_{1}$ and $t_{2}$ represent the block generation time and the parent block generation time, respectively. Note that if the parent block generation time is longer, more fee-valuable transactions will likely remain in the mempool after the parent block is mined, allowing the next block to collect these transactions. As the time-fee equation shows, the fee increase rate with respect to the block generation time is higher than the fee increase rate with respect to the parent block generation time. This implies that the block generation time is a more effective predictor of the transaction fee compared to the parent block generation time.
(a) May 3, 2023
(b) March 18, 2024
Figure 11. Time-fee equation
Appendix E A3C Implementation
E.1. A3C Actions and State Representation
In our A3C-based implementation, we assume that adversary $\mathcal{A}$ mines on top of its secret chain, while the honest miners $\mathcal{H}$ mine on top of the single public chain. These two chains share a common subchain, after which they diverge from each other. The common subchain is referred to as the canonical chain, whose included blocks cannot later be orphaned by the adversarial chain. The adversary’s chain (honest chain) proposed on top of the canonical chain tip is referred to as the adversarial fork (honest fork). Let $l_{\mathcal{A}}$ ($l_{\mathcal{H}}$) denote the adversarial fork length (the honest fork length). We first explain the set of possible actions that the adversary can take.
override: The action represents that the adversary publishes a sub-fork of its secret fork whose length is one block longer than the honest fork. As a result of override action, the honest miners abandon their current fork and start mining on top of the longer, adversarial fork. This action is feasible if $l_{\mathcal{A}}>l_{\mathcal{H}}$.
match: The action represents that the adversary publishes a sub-fork of its secret fork whose length is equal to the honest fork. Action match results in a race between $2$ same-height forks. To encourage altruistic miners to mine on top of the adversarial fork, the adversary should immediately publish its fork after an honest block is released. By doing so, the adversarial fork can be delivered earlier to some of the altruistic nodes, where the ratio of these nodes is determined based on the adversarial communication capability factor. However, petty-compliant miners do not necessarily choose a fork based solely on its delivery time. To incentivize these miners, the adversarial fork must leave transactions with higher total fees in the mempool compared to the honest fork. This means that the adversarial fork should have a lower total transaction fee than the honest fork to improve its chances of winning the same-height fork race in the presence of petty-compliant miners. Typically, if the adversary has mined the block of height $l$ in its fork earlier than the block of height $l$ in the honest fork, the adversarial fork up to and including length $l$ contains a lower total transaction fee compared to the honest fork up to and including length $l$. This is because, during the time gap between mining the $l^{\text{th}}$ adversarial block and the $l^{\text{th}}$ honest block, several fee-rich transactions may arrive that honest miners can include in their blocks, which are absent from the adversarial fork up to and including block $l$. This action is feasible if $l_{\mathcal{A}}\geq l_{\mathcal{H}}$.
wait: This action represents that the adversary continues mining new blocks on top of its secret fork. This action is always feasible.
$\texttt{adopt}_{i}$: The action $\texttt{adopt}_{i}$, $i\in\{0,1,\ldots,k_{1}\}$ represents that the adversary abandons its adversarial fork and starts mining a new fork on top the longest public chain whose latest $i$ blocks are pruned. In the fixed block reward model, considering only the action $\texttt{adopt}_{0}$—namely adopting the longest chain—is enough as $\texttt{adopt}_{i}$ for $i>0$ cannot lead to a more profitable strategy. However, in the volatile block reward model, a deep block in the honest fork may have an extraordinary amount of transaction fees, incentivizing the adversary to orphan the block and steal its transaction fees. A block with an extraordinary fee can result from including a whale transaction as considered in (Bar-Zur et al., 2022) or from having a long generation time. Action $\texttt{adopt}_{i}$ is feasible if $i\leq l_{\mathcal{H}}$.
$\texttt{undercut}_{\texttt{block}}$: This action represents that the adversary tries to undercut the tip block of the longest public chain denoted by $B^{\text{tip}}_{\mathcal{H}}$. To perform block undercutting, the adversary: i) abandons its fork, ii) adopts $l_{\mathcal{H}}-1$ blocks in the honest fork, namely the entire honest fork except for the last block $B^{\text{tip}}_{\mathcal{H}}$, and iii) adjusts the total transaction fee in its potential next block to be less than the total transaction fee included in $B^{\text{tip}}_{\mathcal{H}}$. If, after taking the action $\texttt{undercut}_{\texttt{block}}$, the next block is mined by the adversary, the adversary’s subsequent action is match. Choosing action match will result in a race between two forks of the same height, differing only in their last block. In this scenario, the altruistic miners will continue mining on top of the honest fork since they received it earlier. However, the petty-compliant miners will switch to mining on top of the adversarial fork as it contains a lower amount of transaction fees. Our implementation considers the $\epsilon$-petty-compliant miners with an adjustable $\epsilon$, where the petty-compliant miners deviate from the honest strategy if doing so allows them to earn at least $\epsilon$ BTC more. If the amount of remaining transaction fee in the transaction pool is relatively low and the tip block $B^{\text{tip}}_{\mathcal{H}}$ includes a relatively large amount of transaction fee, undercutting block $B^{\text{tip}}_{\mathcal{H}}$ may be a logical action for the adversary as it enables the adversary to fill its block with transactions with higher fees. This action is always feasible.
$\texttt{undercut}_{\texttt{fork}}$: This action represents that the adversary tries to undercut the entire honest fork using its fork whose length is $1$ less than the honest fork length. To perform fork undercutting, the adversary adjusts the transaction fee in its potential next block in a way that the total transaction fee in the adversarial fork becomes less than that in the honest fork. For $l_{\mathcal{H}}=1$, this action is the same as $\texttt{undercut}_{\texttt{block}}$. Similar to the scenario explained for action $\texttt{undercut}_{\texttt{block}}$, if the next block is mined and published by the adversary, the petty-compliant miners are incentivized to mine on top of the adversarial fork. Under the presence of petty-compliant miners, action $\texttt{undercut}_{\texttt{fork}}$ increases the probability of orphaning the honest fork when the adversary’s fork is lagging behind in the fork race. This action is feasible if $l_{\mathcal{A}}=l_{\mathcal{H}}-1$.
A state at step $t$ in our implementation represents a race between the honest and adversarial forks and has the following format:
(23) $\begin{split}s_{t}=&\Big{(}l_{\mathcal{A}},l_{\mathcal{H}},\texttt{Latest},\texttt{Match},\texttt{undercut},\\ &BR_{\mathcal{A}}=\{BR_{\mathcal{A}}^{1},BR_{\mathcal{A}}^{2},\ldots,BR_{\mathcal{A}}^{\texttt{maxForkLen}}\},\\ &\texttt{Pool}_{\mathcal{A}}=\{\texttt{Pool}_{\mathcal{A}}^{1},\texttt{Pool}_{\mathcal{A}}^{2},\ldots,\texttt{Pool}_{\mathcal{A}}^{\texttt{maxForkLen}}\},\\ &\texttt{Pool}_{\mathcal{H}}=\{\texttt{Pool}_{\mathcal{H}}^{1},\texttt{Pool}_{\mathcal{H}}^{2},\ldots,\texttt{Pool}_{\mathcal{H}}^{\texttt{maxForkLen}}\},\\ &\texttt{Pool}_{\mathcal{C}}\Big{)}\enspace.\end{split}$
In the state representation above, $l_{\mathcal{A}}$ ($l_{\mathcal{H}}$) denotes the adversarial (honest) fork length. Latest is a boolean variable that represents whether the latest mined block is honest or adversarial. If the last block is honest and $l_{\mathcal{A}}\geq l_{\mathcal{H}}$, the adversary can take action match. Match is a boolean variable that represents whether the same-height fork race resulting from the action match is resolved or not. If after taking action match, the next block is honest, the fork race is resolved. However, if the next block is adversarial, the fork race persists. undercut is a boolean variable that represents whether the next block mined after taking actions $\texttt{undercut}_{\texttt{block}}$ or $\texttt{undercut}_{\texttt{fork}}$ is adversarial or not. If the adversary manages to mine the next block, it should then take action match to publish the undercutting block or fork to the rational miners. $BR_{\mathcal{A}}^{i}$ for $i\in\{1,2,\ldots,\texttt{maxForkLen}\}$ denotes the block reward of the $i^{\text{th}}$ block in the adversarial fork. This block reward includes both the protocol reward and the transaction fees collected from the included transactions. Note that maxForkLen is a variable that determines the maximum length of a fork in our implementation. $\texttt{Pool}_{\mathcal{A}}^{i}$ ($\texttt{Pool}_{\mathcal{H}}^{i}$) for $i\in\{1,2,\ldots,\texttt{maxForkLen}\}$ denotes the mempool statistics w.r.t. the chain whose tip block is the $i^{\texttt{th}}$ adversarial (honest) block. Once a new block is mined, it includes some of the transactions available in the mempool, resulting in the removal of those transactions from the mempool and consequently a modification in the mempool statistics. A state must store the mempool statistics w.r.t. the $i^{\texttt{th}}$ adversarial (honest) block as miners may need to abandon their current forks and continue mining on top of the $i^{\texttt{th}}$ adversarial (honest) block as a result of a specific action. Mempool statistics is an $N^{\text{range}}$-element array that contains the weight of transactions available in each of the satoshi-per-byte ranges $\{\texttt{fee}_{0},\texttt{fee}_{1},\ldots,\texttt{fee}_{N^{\text{range}}-1}\}$. The weight of transactions corresponding to each satoshi-per-byte range $\texttt{fee}_{j}$ increases over time according to the specific weight-time equation defined for the range $\texttt{fee}_{j}$. As mentioned earlier, once a new block is mined, a total weight of 1 virtual megabyte is deducted from the transactions in the highest satoshi-per-byte ranges. However, if the adversary takes one of the actions $\texttt{undercut}_{\texttt{block}}$ or $\texttt{undercut}_{\texttt{fork}}$, the block may be filled with less than 1 virtual megabyte of transactions to leave some valuable transactions in the mempool for petty-compliant miners. $\texttt{Pool}_{\mathcal{A}}^{i}$ ($\texttt{Pool}_{\mathcal{H}}^{i}$) is an array containing $N^{\text{range}}$ elements, where the $j^{\text{th}}$ element represents the weight of total transactions of the mempool that pay $\texttt{fee}_{j-1}$ satoshis per virtual byte. Finally, $\texttt{Pool}_{\mathcal{C}}$ denotes the mempool statistics w.r.t. the canonical chain.
E.2. Undercutting as a Continuous Action
The actions override, adopt, match, and wait can be considered discrete actions. This means that once a new block is mined in the system, the agent decides whether to take or not take these actions. However, the action undercut is more appropriately viewed as a continuous action. When dealing with continuous actions, the agent must decide on a value within a range that represents the intensity or duration of the action. In the case of undercut, the adversary must determine how long to continue undercutting the tip of the canonical chain. After a certain point, the adversary should stop undercutting and resume mining on top of the tip block. The continuous nature of the undercut action arises because, depending on the available transaction fees in the mempool, undercutting may or may not be the best action to take. As discussed in Section A, the undercut action is profitable when the mempool w.r.t. the tip of the chain is relatively empty, leading the undercutter to mine on top of the parent block of the tip to steal the transaction fees from the tip block. Over time, as the transaction fees available in the mempool increase, a point is reached where the fees w.r.t. the tip of the chain are large enough that continuing the undercut action becomes unprofitable. At this point, the agent should abandon undercutting and instead mine on top of the tip block.
In an A3C implementation, the neural network outputs a single scalar value representing the value of the current state, along with a probability distribution over the possible actions. For discrete actions, this probability distribution is a vector where each element represents the probability of selecting a specific action given the current state. For continuous actions, the network outputs the parameters of a probability distribution over the continuous action (Jia and Zhou, 2022). In our implementation, the neural network outputs a vector representing the probability distribution of the actions introduced in Section E.1. Additionally, it outputs two scalars representing the mean and variance of the duration of the undercutting action, which is modeled as a continuous action.
E.3. Implementation of the Undercutting Attack
To analyze the undercutting attack, the possible action set for the A3C agents is limited to the following actions:
  • If the adversary mines a new block on top of the longest chain, it publishes the block for the other miners.
  • Whenever the tip of the longest chain, denoted by block $B_{1}$, is not an adversarial block, the adversary attempts to undercut block $B_{1}$. For a period of $t$ seconds, determined by the training process, the adversary mines on the parent of block $B_{1}$. If no block is mined during this time, the adversary abandons the undercutting attempt and resumes mining on top of block $B_{1}$.
  • If, during the undercutting of block $B_{1}$, the next block $B_{2}$ is mined by honest miners on top of $B_{1}$, the adversary stops undercutting block $B_{1}$ and begins undercutting block $B_{2}$.
  • If, while undercutting block $B_{1}$, the adversary successfully mines a competing block $B_{\mathcal{A}}$, it publishes block $B_{\mathcal{A}}$. If the next block $B_{2}$ is mined on top of $B_{1}$, the undercutting attempt fails, and the adversary tries to undercut block $B_{2}$. If block $B_{2}$ is mined on top of block $B_{\mathcal{A}}$, the attack succeeds.
Note that this set of actions does not result in obtaining the optimal strategy and is only designed to measure the profitability of single-block undercutting under different mempool patterns.
E.4. Implementation Details
Our A3C-based tool is adapted from the code provided in (Kostrikov, 2018). Our implementation consists of two linear layers, each with 256 nodes, followed by an LSTM layer that processes the output from the last linear layer. The model uses a learning rate of $10^{-6}$, a discount factor of $0.99$, and an entropy coefficient of $0.01$. We employ 30 training agents to train the neural network and 2 testing agents to evaluate the results. During the training process, information related to the three best checkpoints is stored, and results are presented after running those checkpoints for $10^{6}$ steps.
Appendix F Objective Function Details
F.1. Nonlinear Optimization Function
The value function of the initializing state $s_{0}$ under the time step reward introduced in equation 10 can be obtained as follows:
(24) $V^{\mathcal{A}}(s_{0};\pi,\rho)=\sum_{t=0}^{\infty}{\gamma^{t}R_{\mathcal{A}}(t;\pi)}-\rho\sum_{t=0}^{\infty}{\gamma^{t}\big{(}N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)\big{)}}\enspace.$
Let $\rho_{1}$ be the time-averaged profit achieved under policy $\pi_{1}$, i.e., $\rho_{1}=O^{\mathcal{A}}(\pi_{1})$. If, for a given policy $\pi_{2}$ and as $\gamma\rightarrow 1$, we have $V^{\mathcal{A}}(s_{0};\pi_{2},\rho_{1})>0$, then we can conclude $\pi_{2}$ dominates $\pi_{1}$. This is because $V^{\mathcal{A}}(s_{0};\pi_{2},\rho_{1})>0$ implies that:
$\frac{\sum_{t=0}^{\infty}{\gamma^{t}R_{\mathcal{A}}(t;\pi_{2})}}{\sum_{t=0}^{\infty}{\gamma^{t}\big{(}N_{\mathcal{A}}(t;\pi_{2})+N_{\mathcal{H}}(t;\pi_{2})\big{)}}}>\rho_{1}\enspace.$
According to equation 9 and considering $\gamma\rightarrow 1$, the inequality above results in $O^{\mathcal{A}}(\pi_{2})>O^{\mathcal{A}}(\pi_{1})$, which indicates that following strategy $\pi_{2}$ is more profitable than $\pi_{1}$. Therefore, if the training agents under step rewards $r_{t}(\rho_{1})$ are properly trained, the resulting policy will achieve a time-averaged profit $\rho_{2}$, where $\rho_{2}\geq\rho_{1}$. If, in the next step, the agents are trained based on the step reward $r_{t}(\rho_{2})$, the time-averaged profit of the resulting policy will be $\rho_{3}$ with $\rho_{3}\geq\rho_{2}$. By continuing this process, the trained policy will gradually converge to the optimal policy that yields the highest time-averaged profit.
F.2. Time Step Duration
In our implementation, the duration of each time step follows an exponential distribution with mean $t^{B}$. The duration of each time step is of great importance, as it affects the amount by which transaction weight increases in the mempool. Before the difficulty adjustment, $t^{B}$ is set to 10 minutes. However, after the difficulty adjustment, the value of block generation time $t^{B}$ needs to be adjusted according to the policy followed by the adversary. To calculate the block generation time after difficulty adjustment under policy $\pi$, we need to keep track of the ratio of the number of canonical blocks to the total number of blocks mined under policy $\pi$. To this end, the neural network outputs another value that aims to estimate the following sum:
(25) $\sum_{t=0}^{T}\gamma^{t}\big{(}N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)\big{)}\enspace.$
The neural network is trained to estimate this sum based on the values of $N_{\mathcal{A}}(t;\pi)$ and $N_{\mathcal{H}}(t;\pi)$ at each time step. With access to the sum of the number of canonical blocks and equation 8, the environment can calculate the block generation time. Note that the policy is only trained to maximize the optimization function defined in equation 9, and this sum is not used in training the policy.
Appendix G Selfish Mining After Difficulty Adjustment
In Figure 12, the percentage increase in time-averaged profit of selfish mining after a difficulty adjustment is depicted as a function of the adversarial communication capability. This figure is depicted based on the following assumptions: i) the protocol reward is set to 00 BTC, ii) the mempool pattern observed on December 18, 2023, between 23:00 and 24:00 CET, has occurred, iii) the adversary’s mining share is equal to $\frac{1}{3}$.
Figure 12. The profit percentage increase of selfish mining after difficulty adjustment.
As shown in Figure 12, under the mempool pattern observed on December 18, 2023, difficulty-adjusted selfish mining is more profitable in the fixed reward model compared to the volatile reward model. The main reason for the higher profitability of selfish mining in the fixed reward model is the concavity of the time-fee function during this period. Figure 13 illustrates the time-fee function obtained under the mempool pattern from December 18, 2023, between 23:00 and 24:00 CET. To generate Figure 13, we simulated 100,000 blocks under this mempool pattern, collecting their generation times and fees. As depicted in Figure 13, the corresponding time-fee function for this period is concave. Ignoring the possibility of whale transaction arrivals, according to Theorem 3.1, the concavity of the time-fee function indicates that the ratio of difficulty-adjusted selfish mining profit to honest mining profit under the volatile model is lower than that under the fixed model.
If a sufficient number of whale transactions exist, the substantial fees from these transactions may compensate for the reduced rewards resulting from the concave time-fee function under difficulty-adjusted selfish mining, potentially increasing the profitability of difficulty-adjusted selfish mining in the volatile model. However, as noted in the WeRLman paper (Bar-Zur et al., 2022), these blocks are extremely rare (typically only a few per epoch) and can only marginally enhance profitability in the volatile reward model.
Figure 13. The concave time-fee function.